NHL mega-superstar Auston Matthews scores on 15% of his shots.
What is the expected waiting time before he misses a shot?
1
Expert's answer
2020-05-25T18:58:38-0400

P(miss)=1-P(score)=1-0.15=0.85

expected waiting time:

E(X)=1/P(miss)=1/0.85=1.176 shots